Using uncertainty principle show that an electron can't stay inside the nucleus . Radius of nucleas =5 ×10^-15

1
Expert's answer
2022-02-24T09:05:23-0500

Diameter of the atomic nucleus is of the order of

10^(-15) m


The maximum uncertainity in the position of electron is 10^(-15)m

Mass of electron

= 9.1x 10^(-31) kg

Now,

∆x .∆p = ( h)/( 4pi)


∆x ( m∆v )= ( h)/( 4pi)


∆v = ( h)/( 4pi) x (1 )/(∆xm),

∆v = 5.80 x10^(10) ms^(-1)

This value is much higher than the velocity of light and hence, it is not possible.